# Invariant Subspace Decomposition

### v1

@@ A.4 Further simulation details: MSPE comparison
Appendix B Extension to non-orthogonal subspaces
@@
In Section 2.1 we have defined an orthogonal partition w.r.t. $(X_{t})_{t\in[n]}$ as a collection $\{\mathcal{S}_{j}\}_{j\in\{1,\dots,q\}}$ of pairwise orthogonal linear subspaces of $\mathbb{R}^{p}$ satisfying (6). Finding the orthogonal subspaces that form the partition means in particular finding a rotation of the original $X$-space such that each subspace is spanned by a subset of the rotated axes, and the coordinates of the projected predictors onto one subspace are uncorrelated with the ones in the remaining subspaces. In this section, we show that orthogonality of the subspaces in the partition is not strictly required to obtain a separation of the true time-varying parameter of the form (7), that is, more general invertible linear transformations can be considered besides rotations. In particular, we briefly present results similar to the ones obtained throughout Section 2 but where the subspaces in the partition are not necessarily orthogonal. To do so, we define a collection of (not necessarily orthogonal) linear subspaces $\mathcal{S}_{1},\ldots,\mathcal{S}_{q}\subseteq\mathbb{R}^{p}$ with $\bigoplus_{j=1}^{q}\mathcal{S}_{j}=\mathbb{R}^{p}$ and satisfying (6) a partition w.r.t. $(X_{t})_{t\in[n]}$ (of cardinality $q$), and further call it irreducible if it is of maximal cardinality. A partition w.r.t. $(X_{t})_{t\in[n]}$ can still be identified through a joint block diagonalization of the covariance matrices $(\Sigma_{t})_{t\in[n]}$ as described in Section 2.1.1 but with an adjustment. More specifically, instead of assuming that the joint diagonalizer $U$ is an orthogonal matrix, we only assume it is invertible and for all $j\in\{1,\dots,q\}$, the columns of $U^{S_{j}}$ are orthonormal vectors. A version of Proposition 1 in which the resulting partition is not necessarily orthogonal follows with the same proof. Moreover, similarly to the orthogonal case, the uniqueness of an irreducible partition w.r.t. $(X_{t})_{t\in[n]}$ is implied by the uniqueness of an irreducible non-orthogonal joint block diagonalizer for $(\Sigma_{t})_{t\in[n]}$; explicit conditions under which such uniqueness holds can be found for example in the work by Nion (2011). In the results presented in the remainder of this section we adopt the same notation introduced in Section 2.1.1, and we additionally define the matrix $W\coloneqq U^{-\top}$.
@@
A partition w.r.t. $(X_{t})_{t\in[n]}$ of cardinality $q$ allows us to decompose the true time-varying parameter into the sum of $q$ components. To obtain such a decomposition via non-orthogonal subspaces, oblique projections need to be considered in place of orthogonal ones. Oblique projections are defined (see, e.g., Schott, 2016) for two subspaces $\mathcal{S}_{1},\mathcal{S}_{2}\subseteq\mathbb{R}^{p}$ such that $\mathcal{S}_{1}\bigoplus\mathcal{S}_{2}=\mathbb{R}^{p}$ and a vector $x\in\mathbb{R}^{p}$ as the vectors $x_{1}\in\mathcal{S}_{1}$ and $x_{2}\in\mathcal{S}_{2}$ such that $x=x_{1}+x_{2}$: $x_{1}$ is called the projection of $x$ onto $\mathcal{S}_{1}$ along $\mathcal{S}_{2}$, and $x_{2}$ the projection of $x$ onto $\mathcal{S}_{2}$ along $\mathcal{S}_{1}$. For a partition $\{\mathcal{S}_{j}\}_{j=1}^{q}$ w.r.t. $(X_{t})_{t\in[n]}$, we denote by $P_{\mathcal{S}_{j}\mid\mathcal{S}_{-j}}$ the oblique projection matrix onto $\mathcal{S}_{j}$ along $\bigoplus_{i\in\{1,\dots,q\}\setminus\{j\}}\mathcal{S}_{i}$: this can be expressed in terms of a (non-orthogonal) joint block diagonalizer $U$ corresponding to the partition as $P_{\mathcal{S}_{j}\mid\mathcal{S}_{-j}}=U^{S_{j}}(W^{S_{j}})^{\top}$. Orthogonal partitions $\{\mathcal{S}_{j}\}_{j=1}^{q}$ w.r.t. $(X_{t})_{t\in[n]}$ are a special case of partitions w.r.t. $(X_{t})_{t\in[n]}$. In particular, if the subspaces are pairwise orthogonal, it holds for all $j\in\{1,\dots,q\}$ that $P_{\mathcal{S}_{j}\mid\mathcal{S}^{-j}}=\Pi_{\mathcal{S}_{j}}$.
@@
Similarly to how we define opt-invariance on $[n]$ for orthogonal subspaces in Section 2.1, we say that a subspace $\mathcal{S}_{j}$ in a partition w.r.t. $(X_{t})_{t\in[n]}$ is proj-invariant on $[n]$ if it satisfies for all $t,s\in[n]$ that
@@
By Lemma 4 it follows that for orthogonal partitions proj-invariance is equivalent to opt-invariance. For an irreducible partition w.r.t. $(X_{t})_{t\in[n]}$, we now define the invariant subspace $\mathcal{S}^{\operatorname{inv}}$ and residual subspace $\mathcal{S}^{\operatorname{res}}$ as
@@
It follows directly by the definition of partitions that $\{\mathcal{S}^{\operatorname{inv}},\mathcal{S}^{\operatorname{res}}\}$ is a partition w.r.t. $(X_{t})_{t\in[n]}$. Moreover, $\mathcal{S}^{\operatorname{inv}}$ is proj-invariant on $[n]$ since $P_{\mathcal{S}^{\operatorname{inv}}\mid\mathcal{S}^{\operatorname{res}}}\gamma_{0,t}=\sum_{\begin{subarray}{c}j\in\{1,\dots,q\}:\\ \mathcal{S}_{j}\text{ proj-invariant on }[n]\end{subarray}}P_{\mathcal{S}_{j}\mid\mathcal{S}_{-j}}\gamma_{0,t}$. We finally define the invariant and residual components by
@@
Let $\{\mathcal{S}_{j}\}_{j=1}^{q}$ be a partition w.r.t. $(X_{t})_{t\in[n]}$ and let $U$ be a joint block diagonalizer corresponding to that partition. Then, the following results hold.
@@
Proposition 3 implies in particular that, apart from the joint block diagonalization differences, estimating $\beta^{\operatorname{inv}}$ and $\delta^{\operatorname{res}}_{t}$ in the non-orthogonal case can be done as described in Section 4. Moreover, it holds that the parameter $\beta^{\operatorname{inv}}$ defined for non-orthogonal partitions w.r.t. $(X_{t})_{t\in[n]}$ is still a time-invariant parameter. Under a generalization assumption analogous to Assumption 2, $\beta^{\operatorname{inv}}$ has positive explained variance at all time points $t\in\mathbb{N}$ and can be used to at least partially predict $\gamma_{0,t}$ (we have not added this result explicitly). In addition, also in the non-orthogonal case the estimation of the residual component only requires to estimate a reduced number of parameters, that is, $\operatorname{dim}(\mathcal{S}^{\operatorname{res}})$
@@
These matrices are in block diagonal form and in particular the $2\times 2$ submatrices forming the first diagonal block do not commute. This implies that an irreducible (orthogonal) joint block diagonalizer for these matrices is the three-dimensional identity matrix $I_{3}$, and the diagonal blocks cannot be further reduced into smaller blocks using an orthogonal transformation. It also implies that an irreducible orthogonal partition w.r.t. $(X_{t})_{t\in[n]}$ is given by $\{\mathcal{S}_{j}\}_{j=1}^{2}$ with $\mathcal{S}_{1}\coloneqq\langle e_{1},e_{2}\rangle$ and $\mathcal{S}_{2}=\langle e_{3}\rangle$, where $e_{j}$ is the $j$-th vector of the canonical basis for $\mathbb{R}^{3}$.
@@
$U$ induces an irreducible partition w.r.t. $(X_{t})_{t\in[n]}$ by
@@
Let $\mathcal{S}^{\operatorname{inv}}$ and $\bar{\mathcal{S}}^{\operatorname{inv}}$ be the invariant subspaces associated with the irreducible orthogonal partition $\{\mathcal{S}_{j}\}_{j=1}^{3}$ and the irreducible partition $\{\bar{\mathcal{S}}_{j}\}_{j=1}^{3}$, respectively. As any irreducible orthogonal partition w.r.t. $(X_{t})_{t\in[n]}$ is also a partition w.r.t. $(X_{t})_{t\in[n]}$, it in general holds that
@@
Appendix C Auxiliary results

### v2

@@ A.5 Further simulation details: MSPE comparison
Appendix B ISD estimation algorithm
We provide here the pseudocode summarizing the ISD procedure described in Section 4. The algorithm includes the estimation of the intercept, that is, it considers for all $t\in\mathbb{N}$ the model
$Y_{t}=\gamma_{0,t}^{0}+X_{t}^{\top}\gamma_{0,t}+\epsilon_{t}$ (25)
satisfying the assumptions of Setting 1 but with the addition of $\gamma_{0,t}^{0}\in\mathbb{R}$.
Algorithm 2 ISD: estimation
1:observations $(X_{t},Y_{t})_{t\in[n]\cup\mathcal{I}_{\operatorname{ad}}}$, $X_{t^{*}}$, number of windows $K$, $\lambda\in[0,1]$
2:$\hat{\beta}^{\operatorname{inv}}$, $\hat{\delta}^{\operatorname{res}}_{t^{*}},\hat{\gamma}_{t^{*}},\hat{\gamma}_{t^{*}}^{0}$
3:$(\mathbf{X}_{k})_{k\in[K]}\leftarrow([X_{\frac{(k-1)n}{K}},\dots,X_{\frac{kn}{K}}]^{\top})_{k\in[K]}$
4:$(\mathbf{Y}_{k})_{k\in[K]}\leftarrow([Y_{\frac{(k-1)n}{K}},\dots,Y_{\frac{kn}{K}}]^{\top})_{k\in[K]}$
5:$(\hat{\Sigma}_{k})_{k\in[K]}\leftarrow\{\widehat{\operatorname{Var}}(\mathbf{X}_{k})\}_{k\in[K]}$
6:$\left\{\begin{bmatrix}\hat{\gamma}^{0}_{k}\\ \hat{\gamma}_{k}\end{bmatrix}\right\}_{k\in[K]}\leftarrow\left\{\operatorname{OLS}(\mathbf{Y}_{k},\begin{bmatrix}\mathbf{1}_{\frac{n}{K}}&\mathbf{X}_{k}\end{bmatrix})\right\}_{k\in[K]}$
7:$\begin{bmatrix}\hat{\gamma}^{0}\\ \hat{\gamma}\end{bmatrix}\leftarrow\frac{1}{K}\sum_{k=1}^{K}\begin{bmatrix}\hat{\gamma}^{0}_{k}\\ \hat{\gamma}_{k}\end{bmatrix}$
8:$\hat{U},\{\hat{S}_{j}\}_{j=1}^{q}\leftarrow\operatorname{approxIrreducibleJointBlockDiagonalizer}((\hat{\Sigma}_{t})_{t\in[n]})$triangleright see Sec. 4.1.1
[…]
10:$\hat{S}^{\operatorname{inv}},\hat{S}^{\operatorname{res}}\leftarrow\emptyset$
11:for $j=1,\dots,q$ do
12:     $\Pi_{\hat{\mathcal{S}}_{j}}\leftarrow\hat{U}^{S_{j}}(\hat{U}^{S_{j}})^{\top}$
13:     $(\hat{c}_{k}^{j})_{k\in[K]}\leftarrow(\widehat{\operatorname{Corr}}(\mathbf{Y}_{k}-\mathbf{X}_{k}(\Pi_{\hat{\mathcal{S}}_{j}}\hat{\gamma}),\mathbf{X}_{k}(\Pi_{\hat{\mathcal{S}}_{j}}\hat{\gamma})))_{k\in[K]}$
14:     if $\frac{1}{K}\sum_{k=1}^{K}\left|\hat{c}_{k}^{j}\right|\leq\lambda$  then $\hat{S}^{\operatorname{inv}}\leftarrow\hat{S}^{\operatorname{inv}}\cup\hat{S}_{j}$triangleright see Eq. (19)
[…]
18:$\hat{\beta}^{\operatorname{inv}}\leftarrow\hat{U}^{\operatorname{inv}}\widehat{\operatorname{Var}}(\mathbf{X}\hat{U}^{\operatorname{inv}})^{-1}\widehat{\operatorname{Cov}}(\mathbf{X}\hat{U}^{\operatorname{inv}},\mathbf{Y})$triangleright see Eq. (21)
[…]
20:$\mathbf{X}^{\operatorname{ad}}\leftarrow[(X_{t})_{t\in\mathcal{I}^{\operatorname{ad}}}]^{\top}$, $\mathbf{Y}^{\operatorname{ad}}\leftarrow[(Y_{t})_{t\in\mathcal{I}^{\operatorname{ad}}}]^{\top}$
21:$\hat{\delta}^{\operatorname{res}}_{t^{*}}\leftarrow\hat{U}^{\operatorname{res}}\widehat{\operatorname{Var}}(\mathbf{X}^{\operatorname{ad}}\hat{U}^{\operatorname{res}})^{-1}\widehat{\operatorname{Cov}}(\mathbf{X}^{\operatorname{ad}}\hat{U}^{\operatorname{res}},\mathbf{Y}^{\operatorname{ad}}-\mathbf{X}^{\operatorname{ad}}\hat{\beta}^{\operatorname{inv}})$triangleright see Eq. (22)
[…]
23:if $\{\hat{\gamma}^{0}_{k}\}_{k\in[K]}$ approximately constant then $\hat{\gamma}_{t^{*}}^{0}\leftarrow\hat{\gamma}^{0}$
[…]
25:$\hat{\gamma}_{t^{*}}\leftarrow\hat{\beta}^{\operatorname{inv}}+\hat{\delta}^{\operatorname{res}}_{t^{*}}$triangleright see Eq. (23)
In the estimation of the intercept, the algorithm distinguishes between two cases: (a) the intercept remains approximately constant in $[n]\cup\mathcal{I}_{\operatorname{ad}}$ and (b) the intercept changes with time but is assumed to be approximately constant in $\mathcal{I}_{\operatorname{ad}}$. In case (a), the computation of $\hat{\gamma}_{t^{*}}^{0}$ in line 21 of Algorithm $2$ is done by averaging the estimated values of the (approximately constant) intercept on historical data. Alternatively, in case (a) we could estimate the intercept simultaneously to the invariant component $\beta^{\operatorname{inv}}$: the whole vector (including the intercept) can be computed by taking the OLS solution of regressing $\mathbf{Y}$ on $\begin{bmatrix}\mathbf{1}_{n}&\mathbf{X}\hat{U}^{\operatorname{inv}}\end{bmatrix}$ and premultiplying it by $\begin{bmatrix}\mathbf{e}_{1}&\hat{U}^{\operatorname{inv}}\end{bmatrix}$ (where $\mathbf{e}_{1}$ is the $p$-dimensional vector with the first entry equal to $1$ and the remaining equal to zero) in place of lines 16 and 21 of Algorithm 2 (see Remark 4). In case (b), the intercept can instead be estimated simultaneously to the residual component $\hat{\delta}^{\operatorname{res}}_{t}$: the computation in line 22 of the algorithm is equivalent to taking the first component of the OLS solution of regressing $\mathbf{Y}^{\operatorname{ad}}-\mathbf{X}^{\operatorname{ad}}\hat{\beta}^{\operatorname{inv}}$ on $\begin{bmatrix}\mathbf{1}_{m}&\mathbf{X}^{\operatorname{ad}}\hat{U}^{\operatorname{res}}\end{bmatrix}$, premultiplied by $\begin{bmatrix}\mathbf{e}_{1}&\hat{U}^{\operatorname{res}}\end{bmatrix}$.
Remark 4.
Recall that the population OLS solution for the linear model (25) at time $t$ can be found by adding a $1$ to the vector $X_{t}$ and solving
$\operatorname*{arg\,min}_{\gamma^{0},\gamma}\mathbb{E}\left[Y_{t}-[1\;X_{t}^{\top}]\begin{bmatrix}\gamma^{0}\\ \gamma\end{bmatrix}\right]^{2}.$
The problem has closed form solution
$\begin{bmatrix}\gamma^{0}\\ \gamma\end{bmatrix}=\begin{bmatrix}1&\mathbb{E}[X_{t}]^{\top}\\ \mathbb{E}[X_{t}]&\mathbb{E}[X_{t}X_{t}^{\top}]\end{bmatrix}^{-1}\begin{bmatrix}\mathbb{E}[Y_{t}]\\ \mathbb{E}[X_{t}Y_{t}]\end{bmatrix}=\begin{bmatrix}\mathbb{E}[Y_{t}]-\mathbb{E}[X_{t}]^{\top}\operatorname{Var}(X_{t})^{-1}\operatorname{Cov}(X_{t},Y_{t})\\ \operatorname{Var}(X_{t})^{-1}\operatorname{Cov}(X_{t},Y_{t})\end{bmatrix}.$
Appendix C Extension to non-orthogonal subspaces
@@
In Section 2.1 we have defined an orthogonal and $(X_{t})_{t\in[n]}$-decorrelating partition as a collection $\{\mathcal{S}_{j}\}_{j\in\{1,\dots,q\}}$ of pairwise orthogonal linear subspaces of $\mathbb{R}^{p}$ satisfying (6). Finding the orthogonal subspaces that form the partition means in particular finding a rotation of the original $X$-space such that each subspace is spanned by a subset of the rotated axes, and the coordinates of the projected predictors onto one subspace are uncorrelated with the ones in the remaining subspaces. In this section, we show that orthogonality of the subspaces in the partition is not strictly required to obtain a separation of the true time-varying parameter of the form (7), that is, more general invertible linear transformations can be considered besides rotations. In particular, we briefly present results similar to the ones obtained throughout Section 2 but where the subspaces in the partition are not necessarily orthogonal. To do so, we define a collection of (not necessarily orthogonal) linear subspaces $\mathcal{S}_{1},\ldots,\mathcal{S}_{q}\subseteq\mathbb{R}^{p}$ with $\bigoplus_{j=1}^{q}\mathcal{S}_{j}=\mathbb{R}^{p}$ and satisfying (6) a $(X_{t})_{t\in[n]}$-decorrelating partition (of cardinality $q$), and further call it irreducible if it is of maximal cardinality. A $(X_{t})_{t\in[n]}$-decorrelating partition can still be identified through a joint block diagonalization of the covariance matrices $(\Sigma_{t})_{t\in[n]}$ as described in Section 2.1.1 but with an adjustment. More specifically, instead of assuming that the joint diagonalizer $U$ is an orthogonal matrix, we only assume it is invertible and for all $j\in\{1,\dots,q\}$, the columns of $U^{S_{j}}$ are orthonormal vectors. A version of Proposition 1 in which the resulting partition is not necessarily orthogonal follows with the same proof. Moreover, similarly to the orthogonal case, the uniqueness of an irreducible $(X_{t})_{t\in[n]}$-decorrelating partition is implied by the uniqueness of an irreducible non-orthogonal joint block diagonalizer for $(\Sigma_{t})_{t\in[n]}$; explicit conditions under which such uniqueness holds can be found for example in the work by Nion (2011). In the results presented in the remainder of this section we adopt the same notation introduced in Section 2.1.1, and we additionally define the matrix $W\coloneqq U^{-\top}$.
@@
A $(X_{t})_{t\in[n]}$-decorrelating partition of cardinality $q$ allows us to decompose the true time-varying parameter into the sum of $q$ components. To obtain such a decomposition via non-orthogonal subspaces, oblique projections need to be considered in place of orthogonal ones. Oblique projections are defined (see, e.g., Schott, 2016) for two subspaces $\mathcal{S}_{1},\mathcal{S}_{2}\subseteq\mathbb{R}^{p}$ such that $\mathcal{S}_{1}\bigoplus\mathcal{S}_{2}=\mathbb{R}^{p}$ and a vector $x\in\mathbb{R}^{p}$ as the vectors $x_{1}\in\mathcal{S}_{1}$ and $x_{2}\in\mathcal{S}_{2}$ such that $x=x_{1}+x_{2}$: $x_{1}$ is called the projection of $x$ onto $\mathcal{S}_{1}$ along $\mathcal{S}_{2}$, and $x_{2}$ the projection of $x$ onto $\mathcal{S}_{2}$ along $\mathcal{S}_{1}$. For a $(X_{t})_{t\in[n]}$-decorrelating partition $\{\mathcal{S}_{j}\}_{j=1}^{q}$, we denote by $P_{\mathcal{S}_{j}\mid\mathcal{S}_{-j}}$ the oblique projection matrix onto $\mathcal{S}_{j}$ along $\bigoplus_{i\in\{1,\dots,q\}\setminus\{j\}}\mathcal{S}_{i}$: this can be expressed in terms of a (non-orthogonal) joint block diagonalizer $U$ corresponding to the partition as $P_{\mathcal{S}_{j}\mid\mathcal{S}_{-j}}=U^{S_{j}}(W^{S_{j}})^{\top}$. Orthogonal and $(X_{t})_{t\in[n]}$-decorrelating partitions $\{\mathcal{S}_{j}\}_{j=1}^{q}$ are a special case of $(X_{t})_{t\in[n]}$-decorrelating partitions. In particular, if the subspaces are pairwise orthogonal, it holds for all $j\in\{1,\dots,q\}$ that $P_{\mathcal{S}_{j}\mid\mathcal{S}^{-j}}=\Pi_{\mathcal{S}_{j}}$.
@@
Similarly to how we define opt-invariance on $[n]$ for orthogonal subspaces in Section 2.1, we say that a subspace $\mathcal{S}_{j}$ in a $(X_{t})_{t\in[n]}$-decorrelating partition is proj-invariant on $[n]$ if it satisfies for all $t,s\in[n]$ that
@@
By Lemma 1 it follows that for orthogonal partitions proj-invariance is equivalent to opt-invariance. For an irreducible $(X_{t})_{t\in[n]}$-decorrelating partition, we now define the invariant subspace $\mathcal{S}^{\operatorname{inv}}$ and residual subspace $\mathcal{S}^{\operatorname{res}}$ as
@@
It follows directly by the definition of partitions that $\{\mathcal{S}^{\operatorname{inv}},\mathcal{S}^{\operatorname{res}}\}$ is a $(X_{t})_{t\in[n]}$-decorrelating partition. Moreover, $\mathcal{S}^{\operatorname{inv}}$ is proj-invariant on $[n]$ since $P_{\mathcal{S}^{\operatorname{inv}}\mid\mathcal{S}^{\operatorname{res}}}\gamma_{0,t}=\sum_{\begin{subarray}{c}j\in\{1,\dots,q\}:\\ \mathcal{S}_{j}\text{ proj-invariant on }[n]\end{subarray}}P_{\mathcal{S}_{j}\mid\mathcal{S}_{-j}}\gamma_{0,t}$. We finally define the invariant and residual components by
@@
Let $\{\mathcal{S}_{j}\}_{j=1}^{q}$ be a $(X_{t})_{t\in[n]}$-decorrelating partition and let $U$ be a joint block diagonalizer corresponding to that partition. Then, the following results hold.
@@
Proposition 3 implies in particular that, apart from the joint block diagonalization differences, estimating $\beta^{\operatorname{inv}}$ and $\delta^{\operatorname{res}}_{t}$ in the non-orthogonal case can be done as described in Section 4. Moreover, it holds that the parameter $\beta^{\operatorname{inv}}$ defined for non-orthogonal $(X_{t})_{t\in[n]}$-decorrelating partitions is still a time-invariant parameter. Under a generalization assumption analogous to Assumption 2, $\beta^{\operatorname{inv}}$ has positive explained variance at all time points $t\in\mathbb{N}$ and can be used to at least partially predict $\gamma_{0,t}$ (we have not added this result explicitly). In addition, also in the non-orthogonal case the estimation of the residual component only requires to estimate a reduced number of parameters, that is, $\operatorname{dim}(\mathcal{S}^{\operatorname{res}})$
@@
These matrices are in block diagonal form and in particular the $2\times 2$ submatrices forming the first diagonal block do not commute. This implies that an irreducible (orthogonal) joint block diagonalizer for these matrices is the three-dimensional identity matrix $I_{3}$, and the diagonal blocks cannot be further reduced into smaller blocks using an orthogonal transformation. It also implies that an irreducible orthogonal and $(X_{t})_{t\in[n]}$-decorrelating partition is given by $\{\mathcal{S}_{j}\}_{j=1}^{2}$ with $\mathcal{S}_{1}\coloneqq\langle e_{1},e_{2}\rangle$ and $\mathcal{S}_{2}=\langle e_{3}\rangle$, where $e_{j}$ is the $j$-th vector of the canonical basis for $\mathbb{R}^{3}$.
@@
$U$ induces an irreducible $(X_{t})_{t\in[n]}$-decorrelating partition by
@@
Let $\mathcal{S}^{\operatorname{inv}}$ and $\bar{\mathcal{S}}^{\operatorname{inv}}$ be the invariant subspaces associated with the irreducible orthogonal partition $\{\mathcal{S}_{j}\}_{j=1}^{3}$ and the irreducible partition $\{\bar{\mathcal{S}}_{j}\}_{j=1}^{3}$, respectively. As any irreducible orthogonal and $(X_{t})_{t\in[n]}$-decorrelating partition is also a $(X_{t})_{t\in[n]}$-decorrelating partition, it in general holds that
@@
Appendix D Auxiliary results